OpenStudy (anonymous):

Help please!! Thanks (: Marlene is a sales associate at a department store. Each week she earns $100 plus a 5% commission of all her sales. This week she earned $160. How much, in dollars, were Marlene's sales this week?

OpenStudy (cwrw238):

160 - 100 = $60 60 = 5% so total sales = (60 * 100) / 5 dollars

OpenStudy (cwrw238):

do you follow how i did it? 5% = 5/100

OpenStudy (anonymous):

yes! thank you :)

OpenStudy (cwrw238):

total comes to $1200

OpenStudy (cwrw238):

yw